If the position of the digit of a two digit number are interchanged ,the number obtained is smaller than the original number by 27. If the digits of the number are in the ratio 1:2,what is the original number?
Answer by richwmiller(17219)

x=10a+b,
y=10b+a,
x-y=27,
b/a=1/2
63-36=27
a = 6,
b = 3,
x = 63,
y = 36
63 is the original number.